# An LLM Agent for Functional Bug Detection in Network Protocols

### v1

@@ IV-D RQ3: Efficiency of RFCScan
Setup and Metrics. To quantify the efficiency of RFCScan, we measure input/output token usage, financial cost, and execution time for analyzing each repository. The results are reported separately for Phase 1 (Code Semantic Indexing) and Phase 2 (Retrieval-guided Detection). Although Claude 3.5 is used for both phases in our evaluation, Phase 1 primarily involves lightweight summarization. Since it does not require complex reasoning, it could be replaced by smaller or less expensive LLMs with minimal impact on performance.
@@
Results. As shown in Table V, RFCScan uses an average of 1123K input and 90K output tokens, costs \$5.03, and completes each protocol analysis in 81 minutes. Since the analysis has a complete summarization of the whole code base and compares code against the full RFC, input token usage and cost largely depend on the code base size and RFC length. Overall, RFCScan is efficient for functional bug detection.
@@ IV-E RQ4: Ablation Studies
Total 51.4% 18 8 44.7% 26 9 35.7% 36 8
Ablation Study 1: Without Code Semantic Indexing. To evaluate the impact of code semantic indexing, we disable semantic summarization in Phase 1. The agent still follows the directory structure and explores the codebase hierarchically but no longer uses LLM-generated summaries. Instead, it relies solely on directory and file names plus function signatures. This setup follows the same design of AGENTLESS [23] and LocAgent [22], which uses code structures without semantic summaries. As shown in Table VI, removing semantic indexing significantly reduces precision (81.9% to 51.4%) and results in fewer detected bugs (47 to 26). This highlights the importance of semantic summaries in guiding the agent to more relevant code regions and reducing false positives.
Ablation Study 2: Detection Without Retrieval. To evaluate the impact of the Retrieval stage, we disable all code retrieval tools (Query, Query_Callee, Query_Caller). The detection agent must make decisions using only the initially retrieved function(s), without fetching additional context. As shown in Table VI, this results in a sharp drop in precision (from 81.9% to 44.7%) and fewer detected bugs (from 47 to 35). This shows that retrieval is important for reducing false positives and improving detection quality.
Ablation Study 3: Detection Without Self-critics. To evaluate the impact of self-critics, we disable it during detection. This leads to a higher false positive rate (18.1% to 64.3%) and fewer bugs (47 to 44), emphasizing its crucial role in enhancing detection precision and finding overlooked bugs.
@@ IV-F Case Studies
Bug #16: Incorrect session look up (Demultiplex). As shown in Figure 8(a), the original bfd_find_disc performed an additional check on the peer’s source address after looking up a session by discriminator. This directly contradicts RFC 5880, which mandates that BFD sessions must be demultiplexed only using the Your Discriminator field. Address-based filtering may reject valid sessions when a peer’s source address changes (e.g., due to routing updates or NAT), leading to unintended session loss and network disruption. We submitted a patch to remove the non-compliant address check, and it has been merged by the developers.
Bug #21: LocalDiag not being set (State). In Figure 8(b), the original bfd_set_shutdown function updated the SessionState but failed to assign local_diag. This violated RFC 5880, which requires setting LocalDiag to ADMIN_DOWN when transitioning to the AdminDown state. This may cause peers to misinterpret administrative shutdowns as faults, triggering unnecessary recovery actions and destabilizing the network. We submitted a patch to set local_diag, which has been merged by the developers.
@@
Several threats may affect the validity of our findings. First, we adopt Claude 3.5 Sonnet as our language model with a temperature setting of 0.0 for stable analysis results, thereby enhancing the reproducibility of RFCScan. Nevertheless, alternative models or configurations may yield variations in performance. So we ran controlled experiments using two additional general-purpose API models: GPT-4o and DeepSeek-V3, to analyze the BFD protocol in FRRouting, where RFCScan discovered the most bugs. RFCScan detected 11 and 13 true bugs with the precision of 71.4% and 70.3%, respectively, demonstrating the effectiveness across models. Second, our manual classification of true positives and false positives may introduce subjective bias. To mitigate this risk, two authors independently reviewed each reported functional bug and resolved any disagreements through discussion. We further validated our results by reporting newly discovered bugs to protocol developers, where all bug reports that received a response have been confirmed as true positives. Third, our method relies on high-quality and well-structured documentation such as RFCs. Ambiguous or poorly maintained documentation might hinder semantic understanding, potentially threatening the effectiveness of our technique.
V Related Work
V-A Bug Detection in Network Protocol Implementation
@@
The correctness of network protocol implementations is critical for ensuring secure and reliable digital communication. Existing bug detection techniques for protocol implementations can be broadly categorized into three approaches. First, fuzzing-based techniques—such as BooFuzz [31] and SAGE [32]—primarily identify bugs by triggering crashes during execution. Netlifter [17] and ChatAFL [9] improve fuzzing coverage by using static analysis or LLMs to infer packet formats and generate valid inputs, but still rely on crashes, overlooking subtle semantic issues. LLMIF [33] and mGPTFuzz [10] extract formats or FSMs from specifications using LLMs to uncover semantic bugs, but focus narrowly on state transition or input validation violations and are limited to specific IoT devices, hard to generalize to broader protocol testing. Second, differential analysis techniques [11, 12, 13] detect bugs by comparing the behaviors of multiple independent implementations of the same protocol. These approaches can uncover semantic discrepancies that fuzzing may miss. However, they are inherently limited to scenarios where multiple protocol implementations are available for comparison. Third, formal verification-based techniques rigorously check protocol implementations against formally specified properties. While powerful in reasoning about semantic correctness, they require substantial manual effort to construct, maintain, and validate formal models, posing significant barriers to practical deployment. In contrast to these prior efforts, our proposed technique, RFCScan, addresses functional correctness by bridging the informal specifications in RFC documents with the program semantics of the protocol implementation, facilitating the functional bug detection with high precision and efficiency.
V-B LLM-aided Static Bug Detection
@@
VI Conclusion
This paper presents RFCScan, an LLM agent designed for detecting functional bugs in network protocol implementations by identifying semantic inconsistencies between the code and RFC documents. Specifically, RFCScan comprises two complementary agents: an indexing agent responsible for semantic indexing of source code, and a detection agent performing retrieval-guided inconsistency detection. In our evaluation, RFCScan successfully uncovered 47 functional bugs across six different protocols, of which 20 have been acknowledged or fixed by the original developers. As the first LLM-based approach explicitly targeting functional bug detection in network protocols, our work offers valuable insights into future research on functional bug detection in domain-specific software systems, demonstrating the significant potential of leveraging LLM capabilities in security auditing tasks.

### v2

@@ IV-D RQ3: Efficiency of RFCAudit
Setup and Metrics. To quantify the efficiency of RFCAudit, we measure input/output token usage, financial cost, and execution time for analyzing each repository. The results are reported separately for Phase 1 (Code Semantic Indexing) and Phase 2 (Retrieval-guided Detection). Although Claude 3.5 is used for both phases in our evaluation, Phase 1 primarily involves lightweight summarization. Since it does not require complex reasoning, it could be replaced by smaller or less expensive LLMs with minimal impact on performance.
@@
Results. As shown in Table V, RFCAudit uses an average of 1123K input and 90K output tokens, costs \$5.03, and completes each protocol analysis in 81 minutes. Since the analysis has a complete summarization of the whole code base and compares code against the full RFC, input token usage and cost largely depend on the code base size and RFC length. Overall, RFCAudit is efficient for functional bug detection.
@@ IV-E RQ4: Ablation Studies
Ablation Study 1: Without Code Semantic Indexing. To evaluate the impact of code semantic indexing, we disable semantic summarization in Phase 1. The agent still follows the directory structure and explores the codebase hierarchically but no longer uses LLM-generated summaries. Instead, it relies solely on directory and file names plus function signatures. This setup follows the same design of AGENTLESS [32] and LocAgent [31], which uses code structures without semantic summaries. As shown in Table VI, removing semantic indexing drops precision from 81.9% to 51.4% and reduces detected bugs from 47 to 26. To explain this drop, we analyze each false positive and classify its root cause according to the four categories in Section IV-B3. We observe that false positives due to (i) incomplete context retrieval rise significantly (from 3 to 14), making it the main source of false positives in this setting. Without semantic summaries, the agent confuses similarly named helper functions across modules—for example, retrieving parse_update_subtlv instead of the correct parse_packet when validating Babel update TLVs. This highlights the importance of semantic summarization for accurate context retrieval and reduced false positives.
Ablation Study 2: Detection Without Retrieval. To evaluate the impact of the Retrieval stage, we disable all code retrieval tools (Query, Query_Callee, Query_Caller). The detection agent must make decisions using only the initially retrieved function(s), without fetching additional context. As shown in Table VI, this results in a sharp drop in precision (from 81.9% to 47.1%) and fewer detected bugs (from 47 to 36). This shows that retrieval is important for reducing false positives and improving detection quality. To assess the impact of each retrieval tool, we disable each tool individually. As shown in Table VII, removing any single tool sharply reduces precision from 81.9% with full retrieval to about 47–50%, with little difference across tools. However, the number of detected bugs varies, with the largest drop occurring when Query_Caller is removed (from 47 to 39), suggesting it contributes the most to overall bug coverage.
@@
Total 51.4% 18 8 47.1% 26 10 36.1% 37 8
TABLE VII: Impact of Removing Single Retrieval Tool.
w/o Query w/o Query_Callee w/o Query_Caller
Precision #New #Old Precision #New #Old Precision #New #Old
50.0% 31 10 47.7% 37 8 47.3% 31 8
Ablation Study 3: Detection Without Self-critics. To evaluate the impact of self-critics, we disable it during detection. This leads to a higher false positive rate (18.1% to 63.9%) and fewer bugs (47 to 45), emphasizing its crucial role in enhancing detection precision and finding overlooked bugs.
@@ IV-F Case Studies
We present two representative case studies:
Bug #1: Incorrect request forwarding. RFC 8966 specifies an ordered forwarding rule: when hop count is 2 or more, a node must first attempt to forward via a feasible route (excluding the requester); only if no such route exists should it fall back to a non-feasible one. As shown in Figure 9(a), FRRouting ignored this ordering and always chose a non-feasible route. This misbehavior could cause forwarding loops or prevent requests from reaching valid next hops, undermining routing convergence. We submitted a fix to prioritize feasible routes, which has been merged by the developers.
Bug #35: Miss loop detection. RFC 2332 specifies that if an NHS forwards an NHRP Resolution Reply that lists its own protocol address in the Responder Address Extension, it must discard the packet and generate an NHRP Error Indication of type “Loop Detected”. As shown in Figure 9(b), FRRouting omitted this check, allowing persistent forwarding loops that consume resources and delay address resolution. The bug was confirmed by the developers.
@@
Several threats may affect the validity of our findings. First, we adopt Claude 3.5 Sonnet as our language model with a temperature setting of 0.0 for stable analysis results, thereby enhancing the reproducibility of RFCAudit. Nevertheless, alternative models or configurations may yield variations in performance [40, 41, 42]. So we ran controlled experiments using two additional general-purpose API models: GPT-4o and DeepSeek-V3, to analyze the BFD protocol in FRRouting, where RFCAudit discovered the most bugs. RFCAudit detected 11 and 13 true bugs with the precision of 71.4% and 70.3%, respectively, demonstrating the effectiveness across models. Second, our manual classification of true positives and false positives may introduce subjective bias. To mitigate this risk, two authors independently reviewed each reported functional bug and resolved any disagreements through discussion [41, 43]. We further validated our results by reporting newly discovered bugs to protocol developers, where all bug reports that received a response have been confirmed as true positives. Third, our method relies on high-quality and well-structured documentation such as RFCs. Ambiguous or poorly maintained documentation might hinder semantic understanding, potentially threatening the effectiveness of our technique.
IV-H Evaluation on Industry-Scale Protocol Stacks
We further applied RFCAudit to two large protocol stacks: aws-c-http, a 74K LoC HTTP library in the AWS Common Runtime, and the Linux TCP/IP stack, about 190K LoC in the kernel implementing core transport and network protocols. We evaluated aws-c-http against RFC 7230 (HTTP/1.1), and Linux against RFC 793 (TCP), RFC 768 (UDP), and RFC 791 (IP). On aws-c-http, RFCAudit detected 18 inconsistencies, of which 12 were confirmed as true positives, yielding 12 unique bugs. On the Linux TCP/IP stack, it flagged 3 inconsistencies, with 1 true positive. These results demonstrate that RFCAudit scales to industry-scale codebases and can expose real specification violations with high precision.
V Discussion and Future Work
While our current evaluation targets network protocols, RFCAudit is not limited to this domain. The key requirement is the availability of specification documents that define expected functionality and constraints. In domains such as system libraries, cloud APIs, or security-critical frameworks, specifications are often available as API references, design guidelines, or standards documents. RFCAudit can be adapted to align such specifications with code and flag inconsistencies as potential bugs. We choose network protocols for evaluation because RFCs are detailed and publicly accessible, but extending RFCAudit to other domains is a promising next step to further demonstrate its generality and effectiveness.
VI Related Work
VI-A Bug Detection in Network Protocol Implementation
@@
The correctness of network protocol implementations is critical for ensuring secure and reliable digital communication. Existing bug detection techniques for protocol implementations can be broadly categorized into three approaches. First, fuzzing-based techniques—such as BooFuzz [44] and SAGE [45]—primarily identify bugs by triggering crashes during execution. Netlifter [26] and ChatAFL [15] improve fuzzing coverage by using static analysis or LLMs to infer packet formats and generate valid inputs, but still rely on crashes, overlooking subtle semantic issues. LLMIF [46] and mGPTFuzz [16] extract formats or FSMs from specifications using LLMs to uncover semantic bugs, but focus narrowly on state transition or input validation violations and are limited to specific IoT devices, hard to generalize to broader protocol testing. LTL-Fuzzer [4] uses LTL formulas to guide fuzzing and detect event-ordering violations, but demands significant manual effort for formula crafting and predicate-to-code mapping, and is limited to temporal property violations. Second, differential analysis techniques [18, 19, 20] detect bugs by comparing the behaviors of multiple independent implementations of the same protocol. These approaches can uncover semantic discrepancies that fuzzing may miss. However, they are inherently limited to scenarios where multiple protocol implementations are available for comparison. Third, formal verification-based techniques rigorously check protocol implementations against formally specified properties. While powerful in reasoning about semantic correctness, they require substantial manual effort to construct, maintain, and validate formal models, posing significant barriers to practical deployment. Beyond the three categories, EBugDec [47] targets inconsistencies between RFC evolution and protocol implementations. However, its scope is limited to RFC-evolutionary bugs, primarily focusing on packet parsing issues. In contrast to these prior efforts, our proposed technique, RFCAudit, addresses functional correctness by bridging the informal specifications in RFC documents with the program semantics of the protocol implementation, facilitating the functional bug detection with high precision and efficiency.
VI-B LLM-aided Static Bug Detection
@@
VII Conclusion
This paper presents RFCAudit, an LLM agent designed for detecting functional bugs in network protocol implementations by identifying semantic inconsistencies between the code and RFC documents. Specifically, RFCAudit comprises two complementary agents: an indexing agent responsible for semantic indexing of source code, and a detection agent performing retrieval-guided inconsistency detection. In our evaluation, RFCAudit successfully uncovered 47 functional bugs across six different protocols, of which 20 have been acknowledged or fixed by the original developers. As the first LLM-based approach explicitly targeting functional bug detection in network protocols, our work offers valuable insights into future research on functional bug detection in domain-specific software systems, demonstrating the significant potential of leveraging LLM capabilities in security auditing tasks.
Acknowledgment
We thank all the anonymous reviewers for the insightful and constructive feedback. We are grateful to the Center for AI Safety for providing computational resources. This work was funded in part by the National Science Foundation (NSF) Awards SHF-1901242, SHF-1910300, Proto-OKN 2333736, IIS-2416835, DARPA VSPELLS - HR001120S0058, ONR N00014-23-1-2081, and Amazon. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.